# Intuitive Surgical SurgToolLoc Challenge Results: 2022-2023

### v3

@@ 5 Conclusion
Recent years have produced remarkable advancements in the area of computer vision. This is especially true with many well established public datasets. While the same performance cannot be claimed for similar problems in surgical data science, significant gains are being made. Future work may determine that detection in surgical images is quantifiably more difficult than in other typical image datasets (e.g. COCO, ImageNet). Regardless, we will continue to host this competition since we believe the ability to classify and segment images in these highly unstructured settings is crucial for surgical data science, and the field of computer vision in general.

### v4

@@ 7 Conclusion
Over four years of challenges (2022–2025), working from essentially a single dataset, this series has engaged dozens of teams and highlighted just as many diverse approaches to surgical video-based machine intelligence. In the core task of weakly supervised tool classification and localization, top performances have demonstrated year-over-year progress. The challenges have also broadened in scope. In 2024, we introduced surgical task recognition as a fully supervised category. In 2025, we introduced surgical visual question answering, a multimodal reasoning task that demands integration of visual perception with natural-language generation. The on-going challenge has demonstrated both the broad applicability of our dataset, and the formidability of the machine learning problems that rely on it.
Despite the many skilled solutions employed, team performance on the main challenge has been modest. Weakly supervised tool classification and localization has remained at or below a mAP of 0.5. This is substantially behind fully supervised state-of-the-art methods on general computer vision benchmarks. It is noteworthy that year-over-year improvements have been driven primarily by increasingly sophisticated data curation strategies—human-in-the-loop annotation, pseudo-label generation via tracking and teacher models, and self-supervised pre-training. While the fully-supervised category detection challenge found greater success, the VQA task remains nascent, with BLEU scores that leave considerable room for improvement. Overall this corroborates the inherent difficulty of visual scene understanding in the surgical data science arena and suggests there is room for architectural innovation.
Recent years have produced remarkable advancements in computer vision, particularly with well-established public datasets. While the same performance cannot yet be claimed for surgical data science, the trajectory is encouraging. We will continue to host this competition and evolve its tasks, as we believe that classifying, detecting, and reasoning about surgical scenes in these highly unstructured settings is crucial for the future of surgical data science and computer vision in general. The publicly released dataset accompanying these challenges [1] will, we hope, serve as a lasting resource for the community.